OpenStudy (anonymous):

A homeowner wants several small dormers built on his roof to admit light. They will be in the shape of isosceles triangles whose sides will exceed their base by 5 inches. If the perimeter is 100 inches, what are the lengths of the three sides? The base is a0, and the equal sides are.

OpenStudy (anonymous):

Alright so perimeter is just all sides added together. The base is a and the sides can be expressed as a+5

OpenStudy (anonymous):

So a+a+5+a+5=100

OpenStudy (anonymous):

Can you solve for a?

OpenStudy (anonymous):

confuse

OpenStudy (anonymous):

Alright the base is "a" and the sides are 5 inches longer than the base. So thats the same as a+5

OpenStudy (anonymous):

so it will be a+5+a+5+a+5

OpenStudy (anonymous):

4a+5=100

OpenStudy (anonymous):

Wait

OpenStudy (anonymous):

Take the last 5 off of what you wrote. Let me draw a picture so maybe you can see what im talking about. Give me a second

OpenStudy (anonymous):



OpenStudy (anonymous):

Because both of those sides are 5 inches longer than the base, which is a

OpenStudy (anonymous):

So it would be a+a+5+a+5=100

OpenStudy (anonymous):

ok i see now

OpenStudy (anonymous):

Alright now solve for a

OpenStudy (anonymous):

so what i do now

OpenStudy (anonymous):

3a+10=100

OpenStudy (anonymous):

Yep now subtract 10 from both sides

OpenStudy (anonymous):

And then divide both sides by 3